### v1

# Beyond Flatland: A Geometric Take on Matching Methods for Treatment Effect Estimation

Abstract
Matching is a popular approach in causal inference to estimate treatment effects by pairing treated and control units that are most similar in terms of their covariate information. However, classic matching methods completely ignore the geometry of the data manifold, which is crucial to define a meaningful distance for matching, and struggle when covariates are noisy and high-dimensional. In this work, we propose GeoMatching, a matching method to estimate treatment effects that takes into account the intrinsic data geometry induced by existing causal mechanisms among the confounding variables. First, we learn a low-dimensional, latent Riemannian manifold that accounts for uncertainty and geometry of the original input data. Second, we estimate treatment effects via matching in the latent space based on the learned latent Riemannian metric. We provide theoretical insights and empirical results in synthetic and real-world scenarios, demonstrating that GeoMatching yields more effective treatment effect estimators, even as we increase input dimensionality, in the presence of outliers, or in semi-supervised scenarios.
1 Introduction
Causal inference based on observational data plays a pivotal role in unveiling cause-effect relationships, thereby guiding evidence-based practices across numerous sectors including medicine (Rosenbaum, 2012), public policy (Ben-Michael et al., 2023), epidemiology (Westreich et al., 2017), and econometrics (Sekhon and Grieve, 2012). The goal is to estimate the causal effect of a treatment $T$ on a given outcome $Y$ in the presence of (pre-treatment) input covariates or confounders $X$, as depicted in Figure 1A.
Confounders are responsible for the so-called confounding bias (Rubin, 2005), a fundamental issue in observational studies where there exists an imbalance/discrepancy in the distributions of treated and control units, due to the treatment depending on them instead of being randomly assigned to the observed population. Matching is a well-established approach in causal inference to alleviate confounding bias (Stuart, 2010), where the goal is to pair suitable control units to each treatment unit (and viceversa) based on the similarity of their covariate information.
Traditional matching methods suffer from two main limitations: first, they operate in the space of raw input covariates $\mathcal{X}$ which are noisy and high-dimensional, so it becomes harder to build exact matches (Luo and Zhu, 2017). Indeed, distances become increasingly meaningless for higher dimensions (Aggarwal et al., 2001; Abadie and Imbens, 2006). Second, observations are assumed to live in Euclidean spaces, overlooking the data manifold’s geometry: this may lead to distorted distances Tosi et al. (2014), which result in undesired matched units and imprecise estimates of treatment effects (Stuart, 2010).
In this work, we assume the confounders lie on a manifold of much smaller dimensionality, this is the so called manifold hypothesis. Such manifold structure arises from underlying causal mechanisms that induce statistical correlations among covariates in a constrained manner (Dominguez-Olmedo et al., 2023), as illustrated in Figure 1B. For example, a sport trainer might want to estimate the effect of a particular training strategy (treatment) on athletes performance (outcome). To estimate such effect using observational data, the trainer might need to adjust for a wide variety of covariates, including height, weight and physiological measurements (confounders). Here, two athletes should be more or less similar according to natural differences in their physiology’s: ideally, distances should respect the manifold structure – here, anatomical constraints – such that confounding bias in treatment effect estimation can be reduced.
We propose GeoMatching, a framework to estimate treatment effects accounting for the underlying geometry and uncertainty of the data. Instead of working in the space of raw input characteristics, we learn a low-dim latent representation together with a latent Riemannian metric (a generalization of the Euclidean metric) such that raw distances along the original manifold embedded in the high-dim input space are preserved in the low-dim latent space (see Figure 1C). We can then leverage those geometry-aware distances for matching to estimate treatment effects more accurately, as shown in Figure 1D. The latent representation is a Riemannian manifold, where distance between two data points is given by minimum path length along the data manifold. To our knowledge, this work is the first one looking at geometry-aware methods (in a differential geometric sense) for causal inference.
Figure 1: Key idea of the paper. We propose GeoMatching, a geometry-aware matching framework that pairs cases and controls along the manifold of confounders $X$. A) Assumed causal graph in this work; B) Different sub-causal graphs for covariates $X$ often induce different geometric data structures; latent variable $R$ represents the manifold structure; C) GeoMatching learns a latent Riemannian metric $\mathbf{G}$ (represented as colorbar) which captures the geometric structure and uncertainty of the data, enabling matched pairs to be faithful to the manifold structure. D) Geometry-aware matched samples translate into more accurate estimates of individual (ITE) and average treatment effects (ATE).
Contributions
The contributions of this work are the following: first, we introduce GeoMatching, a geometry-aware matching method that captures the intrinsic data structure of the confounders leveraging Riemannian geometry, accounting for data uncertainty with particular robustness to outliers. We describe assumptions under which we expect GeoMatching to improve over standard matching methods. Second, we empirically show that GeoMatching yields better results compared to other baselines on diverse synthetic and real-world datasets. We show its effectiveness as we increase the input space dimensionality, in the presence of outliers, or in semi-supervised scenarios.
2 Standard Matching and Problem Description
Notation.
Let $(X,T,Y)$ denote input (pre-treatment) covariates $X\in\mathcal{X}$, treatment variable $T\in\{0,1\}$ and outcome variable $Y\in\mathcal{Y}$. We operate in the standard potential outcome framework introduced by Rubin (2005). That is, we assume that each observation has two potential outcomes $Y(0)$ and $Y(1)$ of which only one is observed, $Y=(1-T)Y(0)+TY(1)$, which is the outcome corresponding to the administered treatment $T$. Concretely, imagine that a single individual with pre-treatment covariates $X_{\star}\in\mathbb{R}^{d}$ is exposed to a treatment $T_{\star}=1$, and we observe their outcome $Y_{\star}(1)\in\mathbb{R}$; this is called the treated unit. We also assume access to covariates $X_{j}$ and outcomes $Y_{j}(0)\in\mathbb{R}$ of $m$ individuals not exposed to treatment ($T_{j}=0$) – the pool of available control units. Let $\mu_{t}(x)=\mathbb{E}[Y(t){\,|\,}X=x]$ denote the expected outcome under treatment $T=t$. We make the following standard assumptions from causal inference (Pearl, 2009):
  1. i)
    ignorability: $Y_{i}(1),Y_{i}(0)\perp T_{i}{\,|\,}X_{i}$ – there are no hidden confounders.
  2. ii)
[…]
  3. iii)
    overlap: $\forall x,0<P(T_{i}=1{\,|\,}X_{i}=x)<1$ – there is a chance of receiving either the treatment or control for every unit.
Problem Formulation.
The main goal of matching is to infer the effect of the treatment by comparing the treated outcome $Y_{\star}(1)$ to that of a matched unit $\hat{Y}^{\text{match}}_{\star}$ that plays the role of a hypothetical “twin” of the treated unit had they not been exposed to treatment. In other words, $\hat{Y}^{\text{match}}_{\star}$ aims to be as close as possible to their (unobserved) potential outcome $Y_{\star}(0)$. More formally, we want the matched unit to approximate the expected outcome without treatment, $\mu_{0}(x)=\mathbb{E}[Y(0){\,|\,}X=x]$, i.e., $\mathbb{E}[\hat{Y}^{\text{match}}_{\star}]\approx\mu_{0}(X_{\star})$. A matched unit is constructed as an average of the outcomes of a few control units. In the following and without loss of generality, we focus on the nearest neighbor matching, i.e., the case when a single control unit gets assigned to the treated unit of interest. Let $j^{\dagger}$ denote the index of the nearest neighbor (hypothetical “twin”). A matched unit can then be built as:
$\hat{Y}^{\text{match}}_{\star}=\sum_{j:T_{j}=0}w_{j}Y_{j},\quad\text{where}\quad\begin{cases}w_{j}=1&\text{if }j=j^{\dagger}\text{ (matched)}\\ 0&\text{if }j\neq j^{\dagger}\text{ otherwise}.\end{cases}$ (1)
Note that, if we allow $w_{j}$ to take continuous values, i.e., $0\leq w_{j}\leq 1$, and $\sum_{j}w_{j}=1$, we recover the synthetic controls framework (Abadie, 2021; Curth et al., 2024).
To find the best covariate match $X_{j^{\dagger}}$ for treatment unit $X_{\star}$ we perform nearest neighbours search among all control units according to some distance metric $d(\cdot,\cdot)$, namely:
$j^{\dagger}=\underset{j\in[m]}{\text{argmin}}\;d(X_{\star},X_{j}).$ (2)
The most popular choice is the squared Euclidean distance, i.e., $d(X_{\star},X_{j})=||X_{\star}-X_{j}||^{2}$, which we refer to as standard matching method. However, if there is underlying structure in the confounders given by the natural causal mechanisms of the problem, which is the best distance to use and under which assumptions? This question motivates our contribution GeoMatching, which we describe next.
3 GeoMatching
3.1 Assumptions
Manifold Hypothesis
The manifold hypothesis states that the high-dim confounders $X$ lie near a low-dim (non-linear) manifold embedded in the original input space (Whiteley et al., 2024). Low-dimensional structure arises due to constraints from physical laws and nature, or in other words, geometry often results from the underlying causal mechanisms between different covariates (Dominguez-Olmedo et al., 2023), as illustrated in Figure 1B. The manifold hypothesis is satisfied under common classes of Structural Causal Models; data types where it is often fulfilled include images of 3D objects, phonemes in speech signals, or video streams (Fefferman et al., 2016).
Geometric Faithfulness
We assume that the treatment effect varies smoothly along the manifold of confounders. Back to the motivational example in Section 1, we expect athletes with similar anatomical constraints (reflected by the manifold structure) to respond similarly to different training strategies. As another example, a chemist might want to assess the effect of adding a catalyst to a chemical reaction (treatment) on the chemical reaction rate (outcome) given experimental conditions such as pH, purity of reactants, pressure, temperature, etc (confounders). As we move along the space of possible experimental conditions (manifold), we expect a smooth variation of the treatment effect.
3.2 A Geometric Take on Treatment Effect Matching
The distance between a treated and control unit can be defined as the length of a curve. A Euclidean distance for example corresponds to the length of a straight line. In general, if confounders $X$ lie on a manifold $\mathcal{M}$ (potentially curved space), we define the length of a smooth curve $\gamma:[0,1]\rightarrow\mathcal{M}$ as:
$\mathrm{Length}(\gamma\,;\mathbf{G})=\int_{0}^{1}\sqrt{\gamma^{\prime}(r)^{T}\mathbf{G}(\gamma(r))\gamma^{\prime}(r)dr},$ (3)
where $\gamma^{\prime}(r)=\frac{d}{dr}\gamma(r)$ denotes the velocity of the curve, and $\mathbf{G}$ denotes a Riemannian metric, which is a smooth function that assigns a symmetric positive definite matrix $\mathbf{G}(x)$ to different locations $x\in\mathcal{M}$. Intuitively, $\mathbf{G}$ encapsulates an infinitesimal notion of distance on manifold $\mathcal{M}$.
Standard matching relies on a Riemannian metric $\mathbf{G}=I$, resulting in the squared Euclidean distance:
$d_{\mathcal{E}}(X_{\star},X_{j})=(X_{\star}-X_{j})^{T}\mathbf{I}(X_{\star}-X_{j}).$ (4)
Mahalanobis matching methods rely on a constant Riemannian metric $\mathbf{G}=\Sigma$,
$d_{\mathcal{M}}(X_{\star},X_{j})=(X_{\star}-X_{j})^{T}\boldsymbol{\Sigma}^{-1}(X_{\star}-X_{j}).$ (5)
GeoMatching relies on a Riemannian distance $d_{\mathcal{R}}(X_{\star},X_{j})$, a generalization of the Euclidean and Mahalanobis distance that replaces the constant covariance matrix $\mathbf{I}$ or $\boldsymbol{\Sigma}$ by an input-dependent covariance matrix $\mathbf{G}(x)$, and integrates out the infinitesimal distance contributions along the shortest curve – also called geodesic curve – connecting the two confounders along the manifold.
$d_{\mathcal{R}}(X_{\star},X_{j})=\underset{\gamma_{j}}{\text{min}}\;\mathrm{Length}\Big{(}\gamma_{j}\,;\mathbf{G}\Big{)}.$ (6)
See Appendix 7.1 for further details and precise mathematical definitions of key Riemannian concepts.
3.3 Algorithm Description
We can now introduce GeoMatching, a geometry-aware matching framework that leverages Riemannian distances to pair treated and control units.
Step 1: Learn latent Riemannian manifold.
GeoMatching learns a topology-preserving low-dimensional representation $Z\in\mathcal{Z}$ and latent Riemannian metric $\mathbf{G}$ given confounders $X\in\mathcal{X}$. Considering a latent representation instead of the original input space makes it easier for metric learning, specially if the input space is high-dimensional, and enables to filter out irrelevant information unrelated to the geometry of the confounders. Several methods can be adopted for this stage. One option is to assume a linear projection using Principal Component Analysis (PCA) [^1] Note that PCA is not guaranteed to preserve geometric structure in general, for which other algorithms such as Isomap or Multidimensional Scaling might be more suitable. and fit a parameterized Riemannian metric $\mathbf{G}$. An alternative is to train a probabilistic latent variable model to jointly learn $Z$ and $\mathbb{E}[\mathbf{G}]$, propagating Jacobian information (Tosi et al., 2014). In the experimental Section 5, we resort to the former strategy for simplicity, and posit a parametric Local Inverse Variance (LIV) Riemannian metric (Arvanitidis et al., 2016) in the latent space $\mathcal{Z}$, which is defined as the inverse of a local diagonal covariance matrix with $j$-th entry defined as:
$\mathbf{G}_{jj}(z)=\left(\sum_{i=1}^{N}w_{i}(z)\left(z_{ij}-z_{j}\right)^{2}+\rho\right)^{-1},\quad\text{where}\quad w_{i}(z)=\mathrm{exp}\left(-\frac{||z_{i}-z||^{2}}{2\sigma^{2}}\right).$ (7)
Parameter $\sigma$ controls how fast uncertainty increases as we move away from the manifold; parameter $\rho$ is needed for numerical instability, influencing the magnitude values that the metric can take. LIV assumes a Euclidean metric locally, which is then regularized to have large volume measure in regions of the feature space far away from observations. The inductive bias underlying such metric is that shortest paths on the data manifold should remain close to the observed data, we want curves to avoid crossing low-density high-uncertainty regions. A key advantage of this metric is that geodesic computation is less costly, as off-diagonal terms in metric $\mathbf{G}$ are zero.
Step 2: Compute Riemannian distances.
We compute geodesic or Riemannian distances in the latent representation between treated and control units:
$d_{\mathcal{R}}(X_{\star},X_{j})=\underset{\gamma_{j}}{\text{min}}\;\mathrm{Length}\Big{(}\gamma_{j}\,;\mathbf{G}\Big{)},$ (8)
where $\gamma_{j}:[0,1]\rightarrow\mathcal{Z}$ is a curve connecting the latent representations of the confounders, i.e., $\gamma_{j}(0)=Z_{\star},\,\gamma_{j}(1)=Z_{j}$. To compute geodesic distances, we instantiate a parameterized (discretized) geodesic curve [^2] Considering a discretized curve for the geodesic scales more gracefully than other graph-based approaches, as the discretization of the curve is always one-dimensional independently of the dimension of the latent space. and optimize its parameters by minimizing the following second-order ordinary differential equation (Do Carmo and Flaherty Francis, 1992):
$\gamma^{\prime\prime}=-\frac{1}{2}\mathbf{G}^{-1}\left[\frac{\partial\,\mathrm{vec}\,\mathbf{G}}{\partial\,\gamma}\right]^{T}(\gamma^{\prime}\otimes\gamma^{\prime}),$ (9)
where $\mathrm{vec}\,\mathbf{G}$ stacks the columns of $\mathbf{G}$ and $\otimes$ denotes the Kronecker product. Based on Picard-Lindelöf theorem (Tenenbaum and Pollard, 1985), geodesics are guaranteed to exist and are locally unique given a starting point and initial velocity $\gamma^{\prime}$. See Appendix 7.5 for further details and discussion on computational cost of GeoMatching.
Step 3: Matching along Riemannian manifold.
Given the Riemannian distances from Equation (8), we match treated and control units along the Riemannian manifold by minimizing:
$j^{\dagger}=\underset{j\in[m]}{\operatorname*{arg\,min}}\;d_{\mathcal{R}}(X_{\star},X_{j}).$ (10)
Step 4: Estimate causal treatment effects.
Given the matched units $\hat{Y}^{\text{match}}_{\star}$ defined in Equation (1), we can estimate the individual and average treatment effects as follows:
$\widehat{\mathrm{ITE}}$ $=$ $Y_{\star}-\hat{Y}^{\text{match}}_{\star},$ (11)
$\widehat{\mathrm{ATE}}$ $=$ $\frac{1}{N}\sum^{N}_{i=1}\left(Y_{i}-\hat{Y}^{\text{match}}_{i}\right)^{T_{i}=1}\left(\hat{Y}^{\text{match}}_{i}-Y_{i}\right)^{T_{i}=0}.$ (12)
4 Related Work
Matching in Latent Representations.
Several previous works learn low-dim representations before matching to mitigate undesirable effects of high dimensions in causal inference. Clivio et al. (2022) leverage the representation of intermediate layers of a neural network that predicts the propensity score $p(T|X)$, showing that such representation is approximately a balancing score, i.e., $X\perp T|Z$. Luo and Zhu (2017) and Zhao et al. (2022) propose matching methods on linear projections of the data based on the sufficient dimensionality reduction property, i.e., $X\perp Y|Z$. Wang et al. (2021) and Li et al. (2016) propose matching algorithms for high-dim data: the former relies on variable selection for categorical data; the later conducts and aggregates matching on random projections of the data. Clivio et al. (2023) characterize the bias added to a weighting estimator when using a data representation and propose an objective to learn suitable representations for causal inference. None of these methods account for the geometry of confounders, which is the main goal of this work.
Geometry-aware Machine Learning Models.
Accounting for data geometry has been the object of study of several works in the machine learning community. Tosi et al. (2014) propose to learn a latent representation equipped with a Riemannian metric by assuming a Gaussian Process prior on the mapping function from latents to observations. Arvanitidis et al. (2021), Shao et al. (2017) and Chen et al. (2018) propose to exploit the geometry of the latent representation in deep generative models. These works conclude that treating the latent space as a curved Riemannian space instead of a Euclidean space yields several benefits for clustering, interpolation and prediction. A caveat of those methods is that the computation of geodesics is expensive and selection of hyperparameters for downstream tasks is often non trivial. More recently, Dominguez-Olmedo et al. (2023) show the benefits of accounting for the geometry in the generation of causally-grounded counterfactual explanations for ML classifiers.
Geometry-aware Causal Inference.
While there has been extensive work on learning suitable latent representations for causal inference, very few have considered the geometry of the latent space for treatment effect estimation. Yan et al. (2024) is the only work we found that proposes to consider the geometry of confounders to improve TE estimation. They propose an optimal transport framework to promote the assignment of low weights for outliers, but they still assume Euclidean distances as opposed to a generic Riemannian distance along the data manifold. To our knowledge, our work is the first one to propose geometry-aware treatment effect estimation based on Riemannian geometry.
5 Experiments and Results
We evaluate GeoMatching on synthetic and real-world datasets against other matching strategies in a synthetic swissroll toy scenario, a motion capture data, and two standard causality benchmark datasets. We show that GeoMatching mantains stability as we increase input dimensionality or in the presence of outliers, and benefits from semi-supervised scenarios.
5.1 Experimental Setup
In order to learn the latent Riemannian manifold (Step 1 in Section 3), we project the data using a deterministic dimensionality reduction technique (either PCA or Isomap) and posit a parametric smoothly changing kernel-based Riemannian metric defined by Equation (7). This choice is motivated twofold: i) the resulting pipeline is very simple and easy to implement, with effectively a single parameter $\sigma$ for model selection, ii) we also explored learning the latent Riemannian metric using a Gaussian Process latent variable model, but found that it was less suitable in practice in terms of computational cost and quality of results. In terms of compute resources, we used 1 CPU for each run, with 10 GB and running time of < 24h for each experiment.
Baselines.
We evaluate the performance of GeoMatching according to two dimensions: i) which space gets considered for matching (the original input space $\mathcal{X}$, or a latent space $\mathcal{Z}$), and ii) which distance we rely on (either Euclidean, Mahalanobis, or Riemannian). We also include random matching to have a reference baseline on how difficult it is to match treated and control units on that particular dataset. In terms of latent representation $Z\in\mathcal{Z}$, we consider three alternatives: PCA, Isomap, and the latent representation of Neural Score Matching (NSM) from Clivio et al. (2022) which corresponds to the first layer of a neural network that predicts treatment assignment. Accounting for all combinations of matching space and considered distance, this gives us a total of 9 baselines and 3 ablations of GeoMatching.
Evaluation.
For evaluation, we qualitatively inspect the matched units (geodesic curves) in the latent space, and quantitatively assess performance of treatment effect estimation via Average Treatment Effect (ATE) absolute error, defined as the absolute difference between the true and estimated ATEs. We also report Precision in Estimation of Heterogeneous Effect (PEHE) and distributions of Individual Treatment Effect (ITE) in Appendix 7.3. When plotting the latent representation of GeoMatching, we visualize the magnification factor – also called Riemannian volume – which measures the magnitude of the local distortion of space at location $p\in\mathcal{M}$, i.e., $V_{\mathbf{G}}(p):=\sqrt{\mathrm{det}\,\mathbf{G}(p)}$. Curves that traverse regions with high magnification factor tend to have larger lengths; it can also be seen as how much energy is needed to traverse such space. Geodesics or shortest paths tend to avoid regions with high values of magnification factor, which reflect high uncertainty away from the observed data manifold.
Model Selection.
We use single nearest neighbour matching strategy with replacement for all considered matching methods. We split our data 75/25 into train and test set: we use the train set to optimize parameter $\sigma$ of the LIV Riemannian metric (which controls how fast the metric increases as we move away from the data manifold), by minimizing ATE absolute error in the train set. For the real-world datasets, we also select the dimensionality $K$ of the latent representation based on performance on the train set, sweeping over $K=\{2,3,4,5,6\}$ for Lalonde dataset, and $K=\{2,4,6,10\}$ for IHDP. For synthetic datasets, we fix the latent dimensionality $K=2$. we report performance in both Train and Test set, averaged over 20 training random seeds for synthetic/semi-synthetic datasets, and 5 different training random seeds for real-world scenarios.
5.2 Synthetic Swissroll Dataset
We adapt the classic swissroll dataset (see Appendix 7.2 for precise details on the data generation process) for the task of treatment effect estimation. We generate $N=200$ observations where (pre-treatment) covariates $X$ follow a 3D-swissroll manifold, treatment assignment variables $T$ are Bernoulli draws from a smoothly varying sigmoid function along the swissroll manifold, and outcome variables $Y$ are drawn from linear models. To assess the impact of input dimensionality on the different matching strategies, we embed the three-dimensional swissroll in a space of higher dimensionality $D\in\{3,5,10,15,25,50,100\}$, by adding additional noisy (irrelevant) input dimensions.
For this dataset, we project the input covariates into a 2D latent space. Both PCA and Isomap return similar latent representations, preserving the geometry of the swissroll. Here we report results for PCA. Figure 2 summarizes results on the swissroll dataset. When inspecting matched units (geodesic curves) in the latent space (Figures 2(a) to 2(c)), all baselines disregard the manifold geometry, pairing cases and controls across areas of high uncertainty. Instead, GeoMatching successfully couples cases and controls along the swissroll manifold by accounting for the confounders geometric structure and data uncertainty as we move away from the data manifold.
When looking at performance for treatment effect estimation (Figures 2(d) and 2(e)), GeoMatching yields the most accurate ATEs. As we increase the dimensionality $D$ of input space $\mathcal{X}$, performance of euclidean matching in $\mathcal{X}$ degrades until reaching the same performance as random matching. Unsurprisingly, euclidean matching in $\mathcal{Z}$ remains competitive as $D$ increases since it removes irrelevant dimensions (like GeoMatching) when projecting the data to a low-dim latent space.
(a) Euclidean in $\mathcal{X}$
(b) Euclidean in $\mathcal{Z}$
(c) GeoMatching $\mathcal{Z}$
(d) Treatment Effect Estimation Performance
(e) Impact of input space dimensionality $D$
Figure 2: Results for swissroll synthetic toy dataset. (a-c) Matched pairs according to baselines and our method GeoMatching, (d) ATE Absolute Errors averaged over 20 random seeds.
5.3 Semi-synthetic Human Motion Capture
We use a synthetic scenario of human motion capture data, which consists of real-world input covariates and synthetic treatment and outcome assignment variables. The goal of this experiment is to demonstrate i) effectiveness of GeoMatching handling high-dimensional realistic covariates, and ii) robustness of GeoMatching against outlier confounders, in contrast to geometry-agnostic methods.
Human motion data has often been used to illustrate the impact of data geometry on downstream applications, as latent spaces are expected to be cylindrical or toroidal (doughnut-shaped) (Tosi et al., 2014; Urtasun et al., 2008). Specifically, we consider motion 16 from subject 22 from the CMU Motion Capture Database [^3] http://mocap.cs.cmu.edu which is a repetitive jumping jack motion. Each observation corresponds to a human pose as acquired by a marker-based motion capture system.
We adapt the CMU mocap data for the downstream task of treatment effect estimation. To motivate this new setup, consider Double Dutch, a jump rope game that started in the streets and has now advanced to competitions, since 1974.444https://youtu.be/iiEzf3J4iFk?feature=shared&t=49. To play, two people turn two long ropes in opposite directions while, multiple players jump simultaneously, entering the rope area one at a time. Let $X$ be the motion capture of the current player jumping, $T$ the strategy of entrance, either slow ($T=0$) or fast ($T=1$), and $Y$ the probability of a successful entrance, when ropes do not trip up the new jumper’s feet. To make the data more challenging, we intentionally perturb some time frames to simulate outlier confounders in the dataset (i.e., players jumping in a weird, unnatural manner, or motion capture sensors being defective). See Section 7.2 for further details about the data generation process.
(a) Data Description: CMU Motion Capture Data.
(b) Euclidean in $\mathcal{Z}$
(c) Geodesic in $\mathcal{Z}$
(d) ATE Estimation Performance
Figure 3: Results for semi-synthetic motion capture data. (a) Visualization of the latent space and corresponding high-dimensional confounders. (b-d) Matched pairs according to baselines and our method GeoMatching, (d) Average Treatment Effect Absolute Errors (avg across 20 random seeds).
Figure 3(a) shows a visualization of the data in a 2D-latent space using PCA dimensionality reduction, which preserves the geometric/periodic structure of the data. Figures 3(b) and 3(c) show matched units (geodesic curves) when pairing cases and controls according to Euclidean (baseline) or Riemannian metric (GeoMatching) in the latent space. All baselines are sensitive to the introduced outliers because they do not take into account data uncertainty; in contrast, GeoMatching is robust against the impact of outliers, which results in better ATE estimation, as shown in Figure 3(d). These results align with our a priori expectations, as outliers tend to lie in low-density (high-volume) regions which are unlikely to be traversed by geodesics.
5.4 Real-world Causality Benchmark Datasets
We evaluate GeoMatching on two real-world datasets. We demonstrate that i) GeoMatching is competitive in more realistic scenarios, and ii) GeoMatching can benefit from semi-supervised settings, where we have access to several unlabelled covariate data, from which only a small subset has treatment and outcome information available. We consider the Infant Health Development Program Dataset and the Lalonde [^5] https://users.nber.org/~rdehejia/nswdata2.html dataset, see Appendix 7.2 for further details.
Table 2(b) summarizes the results on real-world datasets. For PCA, GeoMatching improves upon all baselines. We also include results for the Isomap projection. In this case, Euclidean matching in $\mathcal{Z}$ is the best approach, slightly above GeoMatching. We hypothesize that this is due to a coarse grid when optimizing the $\sigma$ parameter of the Riemannian metric (only 10 categorical values were explored in a grid search setup, see Appendix for details). In the case of the LaLonde dataset, using a Riemannian metric yield improvements in the Original covariate space, as well as in the PCA and NSM latent representations.
Table 1: ATE absolute error for real-world datasets.
(a) IHDP dataset.
Original ($\mathcal{X}$) PCA ($\mathcal{Z}$) Isomap ($\mathcal{Z}$) NSM ($\mathcal{Z}$)
Random $1.269\pm 0.053$ NA NA NA
Euclidean 1.126 $\pm$ 0.035 $1.137\pm 0.027$ 1.107 $\pm$ 0.012 1.166 $\pm$ 0.093
Mahalanobis 1.216 $\pm$ 0.063 1.136 $\pm$ 0.042 1.160 $\pm$ 0.051 1.207 $\pm$ 0.060
GeoMatching $1.131\pm 0.036$ 1.118 $\pm$ 0.029 $1.112\pm 0.034$ 1.157 $\pm$ 0.094
(b) Lalonde dataset.
Original ($\mathcal{X}$) PCA ($\mathcal{Z}$) Isomap ($\mathcal{Z}$) NSM ($\mathcal{Z}$)
Random 610.1 $\pm$ 300.5 NA NA NA
Euclidean 541.8 $\pm$ 117.6 364.9 $\pm$ 219.9 1053.2 $\pm$ 921.3 811.5 $\pm$ 591.0
Mahalanobis 809.02 $\pm$ 158.8 843.6 $\pm$ 438.1 766.5 $\pm$ 468.4 948.8 $\pm$ 448.7
GeoMatching 485.6 $\pm$ 319.3 363.2 $\pm$ 122.6 809.6 $\pm$ 456.5 561.4 $\pm$ 396.3
6 Conclusion
In this paper, we adopt a principled differential geometric approach for causal inference. We introduce GeoMatching, a novel matching method for treatment effect estimation that accounts for geometry structure and uncertainty of the confounders. The idea is to match cases and controls based on a latent Riemannian metric instead of a Euclidean metric, such that distances are more meaningful along the data manifold. We learn a non-parametric metric space by constructing a smoothly changing kernel-based deterministic metric that induces a Riemannian manifold. We empirically show that GeoMatching outbeats other baselines for treatment effect estimation, and remains compatitive for increasing input dimensionality, in the presence of outliers, and semi-supervised scenarios.
Limitations
We note that the manifold learning stage, i.e., projecting the data in a way that preserves geometric structure, as well as the computational cost from computing geodesics can be improved. In general, neither PCA nor Isomap are probabilistic approaches, and neither of them are guaranteed to preserve geometric/topological structure. If the GeoMatching is performed in a latent representation that destroys geometric structure, treatment effects could be arbitrarily biased, potentially leading to negative societal impact. This is true for any matching method on latent representations. Computationally, the bottleneck is how to evaluate geodesics, which can be speed-up by minimizing a discretized length of the curve using graph information Chen et al. (2018), or using approximate Riemannian metrics Arvanitidis et al. (2022). We note that GeoMatching is a framework rather than a closed-form procedure, allowing for enhancements in learning better latent representations or accelerating the computation of geodesics to be seamlessly integrated.
Broader Impact
Geometry is an important inductive bias in ML and has been shown to yield benefits in diverse ML applications such as classification, interpolation, model training, and synthetic data generation. To the extend of our knowledge, this work is the first to look into the benefit of geometry-aware methods for causal inference. Estimation of treatment effects using differential geometry is a promising avenue for future research, given that causality often entails geometric manifold structures naturally (Dominguez-Olmedo et al., 2023). As future work, an interesting research direction would be to account for geometry broadly in other causal inference methods, including meta-learners, synthetic controls or propensity-score weighting methods, leveraging Riemannian geometry instead of Euclidean spaces.